# Dense Matter in Neutron Stars with eXTP

### v1

@@ 1 Introduction
The enhanced X-ray Timing and Polarimetry (eXTP) is a science mission primarily designed to study laws of physics under extreme conditions of density (this paper), gravity [1], magnetism [2] and the potential time-domain and multi-messenger transients [3] as well as observatory science [4]. eXTP is currently scheduled to launch in early 2030. In the new baseline design, the scientific payload of eXTP consists of three main instruments: the Spectroscopic Focusing Array (SFA), the Polarimetry Focusing Array (PFA) and the Wide-band and Wide-field Camera (W2C). Here, we provide a brief introduction to the scientific payloads and refer to [5] for a detailed description.
The SFA consists of five SFA-T (where T denotes Timing) X-ray focusing telescopes covering the energy range –, featuring a total effective area of at and at . The designed angular resolution of the SFA is (HPD) with a field of view (FoV). The SFA-T are equipped with silicon-drift detectors (SDDs), which combine good spectral resolution ( 180 eV at 1.5 keV) with very short dead time and a high time resolution of . They are therefore well-suited for studies of X-ray emitting compact objects at the shortest time scales. The SFA array also includes one unit of the SFA-I (where I signifies Imaging) telescope equipped with pn-CCD detectors (p-n junction charged coupled device) to enhance imaging capabilities, which would supply strong upper limits to detect weak and extended sources. The expected FoV of SFA-I is . Therefore, the overall sensitivity of SFA could reach around for an exposure time of . Since it is not excluded that the SFA might in the end include six SFA-T units, simulations presented here have taken this possibility into consideration.
@@
With the large effective area of eXTP, high time resolution, and combined spectral, timing, and polarization capabilities, precise measurements of neutron star (NS) properties hold promise in constraining the NS equation of state (EOS) and the composition of dense matter in its interior [11, 12, 13]. Through the timely efforts enabled by eXTP, we will be able to advance the understanding of several key questions central to the long-standing challenge of determining the EOS of dense matter:
@@ 1.1 State of the art of dense matter
The EOS parameters for asymmetric nuclear matter, such as the symmetry energy or slope of the symmetry energy () around saturation density, can be extracted from experiments involving isospin diffusion measurements and pygmy resonances, isobaric analog states, production of pions and kaons in HIC or data on the neutron skin thickness of heavy nuclei [see reviews e.g., 28, 29, and references therein]. Note that previous nuclear experiments tested matter below or close to saturation density [e.g., 30, 31], whereas HIC probes matter with similar numbers of neutrons and protons up to 2–3 times nuclear saturation density [32, 33] and temperatures well above zero [e.g., 34, 35]. Thus, in order to describe NS matter, models need to be extended to high baryon density in an extremely isospin asymmetric environment at low temperatures [see discussion in e.g., 28, 29, 36, 37]. Nevertheless, it is not obvious whether the information on the nuclear EOS from high-energy HICs can be related to the physics of NS interiors. Methods to integrate diverse data from nuclear experiments and multi-messenger astrophysics to advance the study of dense matter are discussed in Sect. 7.
The regime relevant for NS interiors lies in a parameter space of the phase diagram where QCD is non-perturbative. In contrast, perturbative QCD (pQCD) analysis, thus far, is only reliable at very high densities or temperatures, starting to show signs of convergence at some tens of saturation densities [38, 39]. These calculations provide constraints for the EOS at NS densities due to stability and consistency requirements [40, 41, 42] as discussed below in Sect. 1.3. Perturbative calculations of the EOS have now reached the nearly complete Next-to-Next-to-Next-to-Leading order (N3LO) [43, 39], and the uncertainties arising from missing higher-order terms have been estimated using Bayesian methods [44]. While first-principles lattice QCD [45] can provide convincing information at zero density/chemical potential, it fails at finite density due to the famous sign problem. Lattice data can be extrapolated to finite baryon chemical potential via different systematic schemes, which are, however, only reliable for small and moderate values of with the temperature [46]. In addition, lattice simulations of theories where the sign problem can be avoided, such as QCD at but at finite isospin chemical potential or phase quenched QCD, can provide rigorous constraints to the EOS at large [47, 48].
Using various theories detailed in Sect. 1.3 to describe dense baryonic matter, it has been shown that the conditions inside NSs could allow for the appearance of exotic (non-nucleonic) degrees of freedom, including hyperons, baryons, as well as deconfined quark matter. At large densities in the centers of massive NSs, exotic quark phases such as color superconducting phases have also been predicted. Reviews of EOS physics, including exotic degrees of freedom, can be found in e.g., [49, 50, 51].
Since NSs are very massive, they could also be privileged environments for the accumulation of dark matter [52, 53, 54, 55]. Extensive astronomical evidence supports the existence of dark matter, including galaxy rotation curves, gravitational lensing effects, and precise measurements of the cosmic microwave background radiation. Current studies suggest that dark matter constitutes approximately 27% of the total mass-energy content of the universe, far exceeding the 5% contribution from ordinary baryonic matter. However, despite its crucial role in cosmology, the fundamental nature of dark matter remains largely unknown, and direct detection experiments have so far yielded no definitive results.
NSs serve as unique astrophysical laboratories for investigating dark matter properties. If a significant amount of dark matter exists in NSs, it could profoundly impact various observable properties of NSs, such as the mass-radius relation [56, 53, 57, 54], tidal deformability [58, 59, 60, 61, 62, 63] and evolution [64, 52, 65, 66, 67]. Consequently, these observations can, in turn, be used to constrain the properties of dark matter. In recent years, this topic has attracted considerable attention, with numerous studies exploring various aspects of dark-matter-admixed NSs. It has been found that dark matter can form either a dense core inside the NS or an extended halo surrounding the NS, which would result in different or even opposite effects on some properties of dark-matter-admixed NSs. For example, the formation of a dark halo would increase the gravitational mass and tidal deformability, while a dark core could result in a decrease in the maximum mass, shrink the NS radius and reduce the tidal deformability [see discussion in, e.g., 68].
In the context of X-ray observations, the strong gravitational field of a NS bends the X-rays emitted from its surface. Tracking this deflected light allows for the reconstruction of the external spacetime of the NS, thereby constraining its mass and radius (see discussion in Sect. 2). This technique has been employed by the Neutron star Interior Composition Explorer (NICER) [69] and is expected to play a crucial role in future missions such as eXTP. However, if the NS is surrounded by a dark matter halo, an additional gravitational potential will act on the light emitted from its baryonic surface, leading to modifications in the pulse profile. Studies have shown that these modifications can be significant, reaching up to 10% for a dense dark matter halo [68, 70, 71]. In such cases, measurements based on the assumption of purely baryonic matter may no longer be accurate. To obtain reliable results, both baryonic matter and dark matter must be considered simultaneously in the modeling [72, 73]. Such a strategy may not only allow for a more precise determination of the NS baryonic radius but could also indirectly constrain the mass, interaction strength, and distribution of dark matter particles, providing valuable insights into the fundamental nature of dark matter.
@@ 1.2 Observational properties of pulsars and other NSs
Magnetars are isolated NSs that are largely powered by strong magnetic fields ranging from – G [90]. Magnetars have spin periods of – s and X-ray luminosities as high as [91, 92, 93, 94]. Initially, they were identified as a distinct class of pulsars, with rotational energy typically much less than their X-ray luminosity and relatively slow rotation. This led to the identification of magnetic energy as the main power of both their persistent emission and their different types of bursts [see 95, 96, and references therein]. Observationally, anomalous X-ray pulsars (AXPs) and soft gamma-ray repeaters (SGRs) are now also classified as a single magnetar class. In general, their X-ray spectra can be characterized by one or two thermal components deriving from hot spots on their surface (– keV) and in some cases with one or two power-law components extending to higher energies (in some cases up to keV [97]). The non-thermal component is commonly attributed to resonant cyclotron scattering of the seed surface thermal photons by non-relativistic electrons in the magnetar magnetospheres, or by a relativistic electron component or, in the case of hundred-keV spectra, by a relativistic electron component [98, 99]. Magnetars show a large variety of flaring activity from their long-term outbursts lasting months-years [100], to energetic bursts and flares of magnetospheric origin. The rare and hyper-energetic giant flares from SGRs can have X-ray luminosities up to – with a duration of – s. This distinctive feature may reflect the interaction between the ultra-strong magnetic field and the NS crust or interior.
@@
Accretion-powered pulsars appear when NSs in X-ray binary systems accrete matter from a companion star and their magnetic field is strong enough to channel the material onto the magnetic poles. X-ray binaries can be further classified as high-mass X-ray binaries (HMXBs) or low-mass X-ray binaries (LMXBs), according to the donor star masses. Most X-ray pulsars are observed in HMXBs. Based on the spectral type of the donor star, NS HMXBs are also classified as supergiant X-ray binaries or Be/X-ray binaries (BeXBs) [e.g. 110]. A majority of the HMXB systems are known to be BeXBs with young optical companions of spectral type O or B and they show outburst activity in X-ray [111, 112, 113].
In LMXBs, where the companion has a mass of , the majority of accreting NSs have weak magnetic fields ( G), insufficient to strongly influence accretion, and are non-pulsating, but there are also sources observed as accreting millisecond X-ray pulsars (AMXP). LMXBs may also show recurrent type-I X-ray bursts, which are attributed to unstable thermonuclear burning on the surface of an accreting NS, with typical durations of – s. Most importantly, the spectra associated with the bursts can be used to constrain the properties of NSs [114]. Additionally, the burst light curves may occasionally show pulsations called burst oscillations which are associated with inhomogeneous emission from the surface and can also be used to put constraints on the NS EOS [115].
@@ 1.3.2 The core and uniform nuclear matter
As mentioned in Sect. 1.1, the only regime where fully ab initio calculations for the properties of QCD matter can be performed is that of extremely high densities, beyond those reached in NSs. At lower densities, up to 1–2 times the saturation density, the properties of matter can be calculated from first-principles calculations based on chiral effective field theory (EFT) of QCD [161, 162, 163, 164, 165, 166]. In chiral EFT, the strong interactions between nucleons are given by long-range pion exchanges and short-range contact interactions, which can be organized in a systematic expansion in powers of , where is a typical momentum in the system and - MeV is the breakdown scale [167, 168]. The EFT expansion allows for controlled calculations of atomic nuclei and nucleonic matter including estimates of theoretical uncertainties [169, 170, 171, 172]. In particular, chiral EFT interactions up to N3LO, including two-, three-, and four-nucleon interactions, have been used to calculate the properties of neutron-rich nuclear matter with quantified theoretical uncertainties up to 1–2 times the saturation density in different many-body frameworks [173, 161, 162, 174, 175, 176, 177, 170, 178, 163, 179, 180, 164, 165, 181, 182, 166, 183]. Other model approaches include the use of the Dyson-Schwinger equation [184] and the Nambu–Jona-Lasinio model (NJL) [185, 186, 187]. An extended NJL description of hadronic matter has also been successfully used to describe NS matter [188, 189, 190].
It is known that dense matter within NSs exhibits significant quantum many-body correlation effects in the core regions, such as the strong Landau damping of nucleon quasi-particles and the renormalization of nucleon masses and Fermi velocities [see e.g., 191, 192, 193]. These effects significantly impact the NS EOS and cannot be captured by mean-field or weak-coupling perturbative approaches [194], highlighting the need for non-perturbative methods. The Dyson-Schwinger equation offers a promising framework, allowing unified treatment of the EOS and superfluid transition temperature [195]. Most thermal evolution models (see Sect.3.4) assume a Fermi liquid with a sharp Fermi surface and linear specific heat. In this picture, the direct Urca (dUrca) process is only allowed above a threshold proton fraction [196, 197, 198]. However, in the dense NS core, strong interactions and quantum fluctuations may induce non-Fermi liquid behavior [199], characterized by a blurred Fermi surface and significant many-body corrections to nucleon mass [193]. This can qualitatively alter neutrino emissivity, heat capacity, and thermal conductivity, and may lower the threshold for dUrca processes. Cooling in accreting NS crusts also depends sensitively on nuclear structure inputs [200, 201], while strong magnetic fields modify the electron chemical potential and the number of Landau levels, further affecting Urca rates [202].
@@
Although very general, these approaches have the limitation of being unable to directly determine the internal composition of NSs. In particular, the identification of the presence of a deconfined phase is generally proposed by analyzing the approach to the limit of approximative conformal symmetry, a known property of weakly coupled quark matter that is severely broken in nuclear matter [227, 225]. Since the sound velocity cannot effectively identify the composition of matter at the intermediate densities relevant to compact stars [255], attempts are being made to directly link NS observations and microphysically driven investigations of the EOS across different density regimes, following the description of the RMFT [256, 257], possibly together with a phase transition to deconfined matter within the NJL model [258] or Skyrme energy density functional [259, 260]. This approach supports quantitative studies of the EOS in different density regimes, inferring the values of unresolved nuclear saturation and phase transition properties, especially with the increasing accumulation of robust NS observables. Problems remain in combining the effective interactions with realistic ones to describe the properties of dense matter [261].

### v2

@@ 1 Introduction
The enhanced X-ray Timing and Polarimetry (eXTP) is a science mission primarily designed to study laws of physics under extreme conditions of density (this paper), gravity [1], magnetism [2] and the potential time-domain and multi-messenger transients [3] as well as observatory science [4]. eXTP is currently scheduled to launch in early 2030.
In the new baseline design, the scientific payload of eXTP consists of three main instruments: the Spectroscopic Focusing Array (SFA), the Polarimetry Focusing Array (PFA) and the Wide-band and Wide-field Camera (W2C). The SFA, PFA, and W2C provide complementary observational capabilities on board eXTP to constrain the equation of state (EOS) of dense matter in neutron stars (NSs). We briefly summarize and benchmark their key instrumentation properties—focusing on effective area, background, and polarimetric sensitivity—against other current and planned missions relevant to dense matter research, and refer to [5] for a detailed description. With the large effective area of eXTP, high time resolution, and combined spectral, timing, and polarization capabilities, precise measurements of NS properties hold promise in constraining the NS EOS and the composition of dense matter in its interior [6, 7, 8].
The SFA consists of five SFA-T (where T denotes Timing) X-ray focusing telescopes covering the energy range –, featuring a total effective area of at and at . The designed angular resolution of the SFA is (HPD) with a field of view (FoV). The SFA-T are equipped with silicon-drift detectors (SDDs), which combine good spectral resolution ( 180 eV at 1.5 keV) with very short dead time and a high time resolution of . They are therefore well-suited for studies of X-ray emitting compact objects at the shortest time scales. The SFA array also includes one unit of the SFA-I (where I signifies Imaging) telescope equipped with pn-CCD detectors (p-n junction charged coupled device) to enhance imaging capabilities, which would supply strong upper limits to detect weak and extended sources. The expected FoV of SFA-I is . Therefore, the overall sensitivity of SFA could reach around for an exposure time of —ideal for constraining faint persistent or extended emission. Since it is not excluded that the SFA might in the end include six SFA-T units, simulations presented here have taken this possibility into consideration.
@@
The combined capabilities of eXTP’s three major payloads— the SFA for high-throughput spectroscopy, the PFA for sensitive X-ray polarimetry, and the W2C for transient monitoring —uniquely positions the mission to deliver next-generation constraints on the EOS of NSs across multiple observational messengers and timescales (see Table I in [5] for a benchmark comparison of eXTP’s instrument capabilities against current and planned X-ray missions):
SFA combines high effective area with focusing optics, good spectral resolution (180 eV), and microsecond timing—surpassing the Rossi X-ray Timing Explorer (RXTE) and the Neutron star Interior Composition Explorer (NICER) in soft X-ray sensitivity.
PFA opens a new observational window for dense matter studies by enabling phase-resolved polarimetry of surface hotspots, aiding mass-radius inference via geometric constraints.
W2C offers continuous hard X-ray monitoring (–) over a very wide FoV, enabling detection of rare events like superbursts, long/intermediate X-ray bursts, and glitch-related transients with high cadence and sensitivity.
Through the timely efforts enabled by eXTP, we will be able to advance the understanding of several key questions central to the long-standing challenge of determining the EOS of dense matter:
@@ 1.1 State of the art of dense matter
The regime relevant for NS interiors lies in a region of the QCD phase diagram where the theory is strongly coupled and non-perturbative. First-principles lattice QCD provides robust results at zero baryon chemical potential [20], but fails at finite density due to the sign problem. Extrapolations to finite baryon chemical potential are only reliable at small to moderate [21], though lattice simulations of theories where the sign problem can be avoided, such as QCD at but at finite isospin chemical potential or phase quenched QCD, can provide rigorous constraints to the EOS at large [22, 23].
At extremely high densities, beyond those realized in NSs, perturbative QCD (pQCD) calculations become applicable and provide EOS constraints through stability and asymptotic consistency requirements [24, 25, 26]. These calculations have reached nearly complete Next-to-Next-to-Next-to-Leading order (N3LO) precision [27, 28], with systematic uncertainties from missing higher-order terms quantified via Bayesian methods [29].
At lower densities, up to 1–2 times the saturation density, the properties of matter can be calculated from first-principles calculations based on chiral effective field theory (EFT) of QCD [30, 31, 32, 33, 34, 35]. In chiral EFT, the strong interactions between nucleons are given by long-range pion exchanges and short-range contact interactions, which can be organized in a systematic expansion in powers of , where is a typical momentum in the system and - MeV is the breakdown scale [36, 37]. Up to N3LO, it enables uncertainty-quantified calculations of neutron-rich matter across multiple many-body frameworks [38, 39, 40, 41, 42, 31, 43, 44, 45, 46, 47, 32, 48, 49, 50, 51, 52].
Other model approaches include the use of the Dyson-Schwinger equation [53] and the Nambu–Jona-Lasinio model (NJL) [54, 55, 56]. An extended NJL description of hadronic matter has also been successfully used to describe NS matter [57, 58, 59].
It is known that dense matter within NSs exhibits significant quantum many-body correlation effects in the core regions, such as the strong Landau damping of nucleon quasi-particles and the renormalization of nucleon masses and Fermi velocities [60, 61, 62]. These effects significantly impact the NS EOS and cannot be captured by mean-field or weak-coupling perturbative approaches [63], highlighting the need for non-perturbative methods. The Dyson-Schwinger equation offers a promising framework, allowing unified treatment of the EOS and superfluid transition temperature [64]. Most thermal evolution models (see Sect.3.4) assume a Fermi liquid with a sharp Fermi surface and linear specific heat. In this picture, the direct Urca (dUrca) process is only allowed above a threshold proton fraction [65, 66, 67]. However, in the dense NS core, strong interactions and quantum fluctuations may induce non-Fermi liquid behavior [68], characterized by a blurred Fermi surface and significant many-body corrections to nucleon mass [62]. This can qualitatively alter neutrino emissivity, heat capacity, and thermal conductivity, and may lower the threshold for dUrca processes. Cooling in accreting NS crusts also depends sensitively on nuclear structure inputs [69, 70], while strong magnetic fields modify the electron chemical potential and the number of Landau levels, further affecting Urca rates [71].
@@ 1.2 Observational properties of pulsars and other NSs
Magnetars are isolated NSs that are largely powered by strong magnetic fields ranging from – G [106]. Magnetars have spin periods of – s and X-ray luminosities as high as [107, 108, 109, 110]. Initially, they were identified as a distinct class of pulsars, with rotational energy typically much less than their X-ray luminosity and relatively slow rotation. This led to the identification of magnetic energy as the main power of both their persistent emission and their different types of bursts [see 111, 112, and references therein]. Observationally, anomalous X-ray pulsars (AXPs) and soft gamma-ray repeaters (SGRs) are now also classified as a single magnetar class. In general, their X-ray spectra can be characterized by one or two thermal components deriving from hot spots on their surface (– keV) and in some cases with one or two power-law components extending to higher energies (in some cases up to keV [113]). The non-thermal component is commonly attributed to resonant cyclotron scattering of the seed surface thermal photons by non-relativistic electrons in the magnetar magnetospheres, or by a relativistic electron component or, in the case of hundred-keV spectra, by a relativistic electron component [114, 115]. Magnetars show a large variety of flaring activity from their long-term outbursts lasting months-years [116], to energetic bursts and flares of magnetospheric origin. The rare and hyper-energetic giant flares from SGRs can have X-ray luminosities up to – with a duration of – s. This distinctive feature may reflect the interaction between the ultra-strong magnetic field and the NS crust or interior.
@@
Accretion-powered pulsars appear when NSs in X-ray binary systems accrete matter from a companion star and their magnetic field is strong enough to channel the material onto the magnetic poles. X-ray binaries can be further classified as high-mass X-ray binaries (HMXBs) or low-mass X-ray binaries (LMXBs), according to the donor star masses. Most X-ray pulsars are observed in HMXBs. Based on the spectral type of the donor star, NS HMXBs are also classified as supergiant X-ray binaries or Be/X-ray binaries (BeXBs) [e.g. 126]. A majority of the HMXB systems are known to be BeXBs with young optical companions of spectral type O or B and they show outburst activity in X-ray [127, 128, 129].
In LMXBs, where the companion has a mass of , the majority of accreting NSs have weak magnetic fields ( G), insufficient to strongly influence accretion, and are non-pulsating, but there are also sources observed as accreting millisecond X-ray pulsars (AMXP).
LMXBs may also show recurrent type-I X-ray bursts, which are attributed to unstable thermonuclear burning on the surface of an accreting NS, with typical durations of – s.
Most importantly, the spectra associated with the bursts can be used to constrain the properties of NSs [130]. Additionally, the burst light curves may occasionally show pulsations called burst oscillations which are associated with inhomogeneous emission from the surface and can also be used to put constraints on the NS EOS [131].
@@ 1.3.2 The core and uniform nuclear matter
The EOS parameters for asymmetric nuclear matter, such as the symmetry energy or slope of the symmetry energy () around saturation density, can be extracted from experiments involving isospin diffusion measurements and pygmy resonances, isobaric analog states, production of pions and kaons in HIC or data on the neutron skin thickness of heavy nuclei [see reviews e.g., 185, 175, and references therein]. Note that previous nuclear experiments tested matter below or close to saturation density [e.g., 186, 187], whereas HIC probes matter with similar numbers of neutrons and protons up to 2–3 times nuclear saturation density [188, 189] and temperatures well above zero [e.g., 190, 191]. Thus, in order to describe NS matter, models need to be extended to high baryon density in an extremely isospin asymmetric environment at low temperatures [see discussion in e.g., 185, 175, 192, 193]. Nevertheless, it is not obvious whether the information on the nuclear EOS from high-energy HICs can be related to the physics of NS interiors. Methods to integrate diverse data from nuclear experiments and multi-messenger astrophysics to advance the study of dense matter are discussed in Sect. 7.
Using various theories detailed in Sect. 1.1 to describe dense baryonic matter, it has been shown that the conditions inside NSs could allow for the appearance of exotic (non-nucleonic) degrees of freedom, including hyperons, baryons, as well as deconfined quark matter. At large densities in the centers of massive NSs, exotic quark phases such as color superconducting phases have also been predicted. Reviews of EOS physics, including exotic degrees of freedom, can be found in e.g., [72, 194, 195].
@@
Although very general, these approaches have the limitation of being unable to directly determine the internal composition of NSs. In particular, the identification of the presence of a deconfined phase is generally proposed by analyzing the approach to the limit of approximative conformal symmetry, a known property of weakly coupled quark matter that is severely broken in nuclear matter [202, 200]. Since the sound velocity cannot effectively identify the composition of matter at the intermediate densities relevant to compact stars [230], attempts are being made to directly link NS observations and microphysically driven investigations of the EOS across different density regimes, following the description of the RMFT [231, 232] or Skyrme energy density functional [233, 234], possibly together with a phase transition to deconfined matter within the NJL model [235]. This approach supports quantitative studies of the EOS in different density regimes, inferring the values of unresolved nuclear saturation and phase transition properties, especially with the increasing accumulation of robust NS observables. Problems remain in combining the effective interactions with realistic ones to describe the properties of dense matter [236].
@@
NSs, due to their high densities and strong gravity, are promising astrophysical laboratories for probing the properties of dark matter [247, 248, 249, 250]. Although dark matter constitutes about 27% of the universe’s energy budget—far exceeding the 5% from ordinary matter—its fundamental nature remains unknown, and direct detection efforts have so far been inconclusive. If dark matter accumulates in NSs, it could significantly alter observable properties such as the mass-radius relation [251, 248, 252, 249], tidal deformability [253, 254, 255, 256, 257, 258], and evolutionary processes [259, 247, 260, 261, 262]. Depending on its distribution, dark matter may form a compact core or an extended halo, leading to contrasting effects: a dark matter halo tends to increase mass and deformability, while a dark matter core can reduce the maximum mass and shrink the radius [263]. These signatures offer a valuable avenue for constraining dark matter through NS observations.
In the context of X-ray observations, the strong gravitational field of a NS bends the X-rays emitted from its surface. Tracking this deflected light allows for the reconstruction of the external spacetime of the NS, thereby constraining its mass and radius (see discussion in Sect. 2). This technique has been employed by NICER [264] and is expected to play a crucial role in future missions such as eXTP. However, if the NS is surrounded by a dark matter halo, an additional gravitational potential will act on the light emitted from its baryonic surface, leading to modifications in the pulse profile. Studies have shown that these modifications can be significant, reaching up to 10% for a dense dark matter halo [263, 265, 266]. In such cases, measurements based on the assumption of purely baryonic matter may no longer be accurate. To obtain reliable results, both baryonic matter and dark matter must be considered simultaneously in the modeling [267, 268]. Such a strategy may not only allow for a more precise determination of the NS baryonic radius but could also indirectly constrain the mass, interaction strength, and distribution of dark matter particles, providing valuable insights into the fundamental nature of dark matter.